# Meta-Unlearning on Diffusion Models: Preventing Relearning Unlearned Concepts

### v1

@@ 4 Meta-unlearning for DMs
In Section 3.2, we have briefly introduced the commonly used unlearning methods for DMs. In general, the unlearning objective can be summarized as forgetting knowledge from $\mathcal{D}_{\textrm{forget}}$ and preserving performance on $\mathcal{D}_{\textrm{retain}}$, i.e., solving
$\min_{\theta}\mathcal{L}_{\textrm{unlearn}}\left(\theta;\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{retain}}\right)\triangleq\mathcal{L}_{\textrm{forget}}\left(\theta;\mathcal{D}_{\textrm{forget}}\right)+\lambda\cdot\mathcal{L}_{\textrm{retain}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)\textrm{,}$ (6)
where $\mathcal{L}_{\textrm{forget}}$ is to unlearn the forget set, $\mathcal{L}_{\textrm{retain}}$ is to preserve performance on the retain set, and $\lambda$ is a trade-off hyperparameter. Various unlearning methods correspond to different instantiations of $\mathcal{L}_{\textrm{forget}}$ and $\mathcal{L}_{\textrm{retain}}$. In particular, ESD and SDD require optimizers to solve their instantiations, whereas UCE and RECE have closed-form solutions. To solve Eq. (6), the initialization is usually set to the pretrained parameters $\theta^{*}$, and the unlearned model parameters are denoted as $\theta^{\textrm{UN}}$.
Algorithm 1 The general framework of meta-unlearning
1:Pretrained parameters $\theta^{*}$, forget set $\mathcal{D}_{\text{forget}}$, retain set $\mathcal{D}_{\text{retain}}$
3:Outer (steps $N$, learning rate $\omega$), inner (steps $M$, learning rate $\tau$), scale factors $\gamma_{1},\gamma_{2}$
4:$\theta_{0}\leftarrow\theta^{*}$triangleright If $\mathcal{L}_{\textrm{unlearn}}$ is ESD/SDD that needs optimization
5:$\theta_{0}\leftarrow\theta^{\textrm{UN}}=\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\textrm{unlearn}}$triangleright If $\mathcal{L}_{\textrm{unlearn}}$ is UCE/RECE that has closed-form solution
6:for $n=1$ to $N$ do
7:     Sample a finetuning set $\mathcal{D}_{\text{FT}}\subset\mathcal{D}_{\text{forget}}$
8:     Initialize $\bm{g}=\bm{0}$ and $\theta^{\textrm{FT}}=\theta_{n-1}$
9:     $\bm{g}\leftarrow\bm{g}+\gamma_{1}\cdot\nabla_{\theta_{n-1}}\mathcal{L}_{\text{unlearn}}(\theta_{n-1};\mathcal{D}_{\text{forget}},\mathcal{D}_{\text{retain}})$triangleright If $\mathcal{L}_{\textrm{unlearn}}$ is ESD/SDD
10:     for $m=1$ to $M$ do
11:         $\theta^{\textrm{FT}}\leftarrow\theta^{\textrm{FT}}-\tau\cdot\nabla_{\theta^{\textrm{FT}}}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$
12:     end for
13:     $\bm{g}\leftarrow\bm{g}+\gamma_{2}\cdot\nabla_{\theta_{n-1}}\mathcal{L}_{\text{meta}}(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})$triangleright Meta objective
14:     $\theta_{n}\leftarrow\theta_{n-1}-\omega\cdot\bm{g}$
15:end for
16:return $\theta_{N}$
@@ 4.1 Meta-unlearning framework
A publicly released DM can potentially be finetuned to adapt to various downstream tasks. However, as observed in previous studies, finetuning or modifying weights of a model could comprise its alignment and/or unlearning (Qi et al., 2023; Tamirisa et al., 2024). This underscores the need for mechanisms to simulate the finetuning process in advance, ensuring DMs are resilient against relearning the unlearned concepts. Inspired by meta-learning (Finn et al., 2017), we propose the meta-unlearning framework, as illustrated in Algorithm 1. Our framework consists of two components: (1) the standard unlearning objective $\mathcal{L}_{\textrm{unlearn}}$, as described above, and (2) the meta objective $\mathcal{L}_{\textrm{meta}}$, which resists the relearning of unlearned concepts, even after finetuning on the forget set.
Formally, we define $\mathcal{L}_{\textrm{FT}}$ as the finetuning objective, and let $\mathcal{D}_{\textrm{FT}}\subset\mathcal{D}_{\textrm{forget}}$ represent the malicious finetuning dataset, which is designed to intentionally make the model relearn concepts from the forget set. The finetuned model parameters $\theta^{\textrm{FT}}$ are updated by one or more gradient descents. For example, when using one gradient update from $\theta$, there is $\theta^{\textrm{FT}}\leftarrow\theta-\tau\cdot\nabla_{\theta}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$, where $\tau$ is the step size. The model parameters $\theta$ is trained by minimizing the meta objective $\mathcal{L}_{\textrm{meta}}$ as:
$\min_{\theta}\mathcal{L}_{\textrm{meta}}(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})=\mathcal{L}_{\textrm{meta}}(\theta-\tau\cdot\nabla_{\theta}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right);\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})\textrm{.}$ (7)
To optimize the right hand side of Eq. (7), the gradients are back-propagated through both $\theta$ and $\nabla_{\theta}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ that can be efficiently computed by automatic differentiation (Paszke et al., 2019).
@@ 4.2 Meta objective $\mathcal{L}_{\textrm{meta}}$
$\min_{\theta}\mathcal{L}_{\textrm{meta}}(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})\triangleq-\mathcal{L}_{\textrm{DM}}(\theta^{\textrm{FT}};\mathcal{D}_{\textrm{FT}})-\zeta\cdot\left[\mathcal{L}_{\textrm{DM}}(\theta^{\textrm{FT}};\mathcal{D}_{\textrm{retain}})-\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{retain}})\right]\textrm{,}$ (8)
@@
$\begin{split}&\min_{\theta}\mathcal{L}_{\textrm{meta}}(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})\\ =&-\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{FT}})+\tau\cdot{\color[rgb]{1.0,0.44,0.37}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.44,0.37}\left\|\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)\right\|_{2}^{2}}+\tau\zeta\cdot{\color[rgb]{1.0,0.44,0.37}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.44,0.37}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)}\textrm{,}\end{split}$ (9)
@@
Remark. As illustrated in Fig. 1, minimizing $\left\|\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)\right\|_{2}^{2}$ decreases the finetuning gradient norm and thus delay the relearning of forget set. Minimizing $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$ induces a $>90^{\circ}$ angle between $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ and $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$, such that when th model is finetuned along $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ (the loss $\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ decreases), the knowledge inside the retain set will self-destruct (namely, the loss $\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$ increases).
@@
5 Experiments
We first describe the basic setups of our experiments, which are outlined below:
Base models. We choose SD-v1-4 (Rombach et al., 2022) and SDXL (Podell et al., 2023) as the base models for their widespread use and strong generation capabilities.
Datasets. We use SD-v1-4 and SDXL to generate both $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{retain}}$ for meta-unlearning. Subsequently, we employ FLUX.1 (Labs, 2024) to create three safety-related finetuning datasets HRM-s, HRM-m, CLEAN for evaluation, by applying a single harmful prompt, multiple harmful prompts and benign prompts, respectively. We also utilize FLUX.1 to create two finetuning datasets related to copyright and two others related to image style. Detailed information can be found in Appendix E.1. Additionally, the 10K subset of COCO-30K (Lin et al., 2014) is used to evaluate the generation quality of unlearned DMs while the nudity subset in the Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023) is used to test the unlearning performance.
Figure 3: Images generated by harmful prompts. The top panel displays images generated using the original SDXL model for harmful prompts. In the following panels, we show images generated using unlearned and meta-unlearned SDXL models before finetuning (FT), after FT on the HRM-m dataset for 100 steps, and after FT on the CLEAN dataset for 100 steps, respectively. The left three columns display images generated by ESD-u-1 and its meta-learning variant, while the right three columns display images generated by UCE and its meta-learning variant.
Baselines. We use four established unlearning methods as baselines, including ESD (Gandikota et al., 2023) and SDD (Kim et al., 2023), which remove the target concept through gradual optimization; UCE (Gandikota et al., 2024) and RECE (Gong et al., 2024) that achieve target concept erasure through closed-form solutions. Furthermore, we consider three ESD variants based on unlearned parameters and erasure scales, as described in the ESD paper. We use the ESD-u-1, which erases U-Net models excluding cross-attention parameters under weak erase scale $\eta=1$, ESD-u-3, which erases the same parameters as ESD-u-1 but under strong erase scale $\eta=3$, and ESD-f-3, which erases the full parameters of the U-Net with erase scale $\eta=3$.
Evaluation metrics. We use FID (Heusel et al., 2017) and CLIP scores (Hessel et al., 2021) to evaluate the model’s generation quality. To evaluate each method’s unlearning performance on harmful content and resistance to malicious finetuning, we calculate the nudity score (Schramowski et al., 2023) based on the percentage of nude images in all generated images.
Evaluation details. We first use an unlearned model to generate images based on COCO 10K subset, and compute the FID and CLIP scores using the generated image and COCO subset data. Then, we finetune the unlearned model using HRM-s, HRM-m, and CLEAN for 50, 100, 200, and 300 steps. Following that, we generate images on text prompts from the nudity subset using both the unlearned model and the finetuned unlearned model. Finally, we use the nudity detector (Zhang et al., 2023; Schramowski et al., 2023) to determine the nudity score for the generated images.
5.1 Unsafe content removal
Tables 1 and 2 show the evaluation results of the unlearned and meta-unlearned SD-v1-4. Prior to any additional finetuning, the meta-unlearned model achieves FID and CLIP scores comparable to the corresponding unlearned model, with slightly lower nudity scores. After malicious finetuning on the HRM-s and HRM-m datasets, the unlearned model shows a rapid increase in nudity scores. In contrast, meta-unlearned SD yields significantly lower nudity scores than the unlearned model. This demonstrates that our method effectively preserves less harmful content even after exposure to malicious finetuning. Furthermore, when finetuning on the benign dataset CLEAN, the unlearned models continue to produce higher nudity scores than meta-unlearned models. Fig. 10 shows images generated by unlearned and meta-unlearned models on benign prompts before finetuning, indicating that meta-unlearned models can produce comparable images with corresponding unlearned models. Then we show the images generated on harmful prompts in Fig. 11. The unlearned and meta-unlearned models are finetuned on the HRM-m dataset for 50, 100, 200, and 300 steps. As finetuning steps increase, unlearned models rapidly relearn the ability to generate harmful images. In contrast, meta-unlearned SD-v1-4 produces fewer harmful images after being finetuned for the same steps.
Fig. 8 presents images generated on benign prompts by unlearned and meta-unlearned SDXL. It can be observed that, for benign prompts, the meta-unlearned SDXL also achieves a high generation quality comparable to that of the corresponding unlearned method. Fig. 3 displays the harmful images generated by the unlearned and meta-unlearned models before and after being finetuned on the HRM-m and CLEAN datasets. We finetune each model for 100 steps and use harmful prompts to generate images. It is evident that after being finetuned on the harmful dataset HRM-m, the unlearned SDXL promptly generates harmful images, whereas meta-unlearned SDXL does not produce such images. Furthermore, after being finetuned on the benign dataset CLEAN, the unlearned models still have a probability of generating harmful images, while meta-unlearned models consistently ensure the generation of harmless images.

### v2

@@ 4 Meta-unlearning for DMs
In Section 3.2, we briefly introduce the commonly used unlearning methods for DMs. In general, the unlearning objective can be summarized as forgetting knowledge from $\mathcal{D}_{\textrm{forget}}$ and preserving performance on $\mathcal{D}_{\textrm{retain}}$, i.e., solving
$\begin{split}&\min_{\theta}\mathcal{L}_{\textrm{unlearn}}\left(\theta;\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{retain}}\right)\\ \triangleq&\mathcal{L}_{\textrm{forget}}\left(\theta;\mathcal{D}_{\textrm{forget}}\right)+\lambda\cdot\mathcal{L}_{\textrm{retain}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)\textrm{,}\end{split}$ (1)
where $\mathcal{L}_{\textrm{forget}}$ is to unlearn the forget set, $\mathcal{L}_{\textrm{retain}}$ is to keep performance on the retain set, and $\lambda$ is a trade-off hyperparameter. Various unlearning methods correspond to different instantiations of $\mathcal{L}_{\textrm{forget}}$ and $\mathcal{L}_{\textrm{retain}}$. In particular, ESD and SDD require optimizers to solve Eq. (1), whereas UCE and RECE have closed-form solutions. In optimization, the initialized parameters are usually set to the pretrained $\theta^{*}$ and the unlearned parameters are denoted as $\theta^{\textrm{UN}}$.
@@ 4.1 Meta-unlearning framework
A publicly released DM can potentially be finetuned to adapt to various downstream tasks. However, as observed in previous studies, finetuning or modifying weights of a model could comprise its alignment and/or unlearning [48, 59]. This underscores the need for mechanisms to simulate the finetuning process in advance, ensuring DMs are resilient against relearning the unlearned concepts. Inspired by meta-learning [5], we propose the meta-unlearning framework, as illustrated in Algorithm 1. Our framework consists of two components: (1) the standard unlearning objective $\mathcal{L}_{\textrm{unlearn}}$, as described above, and (2) the meta objective $\mathcal{L}_{\textrm{meta}}$, which resists the relearning of unlearned concepts, even after finetuning on the forget set.
Formally, we define $\mathcal{L}_{\textrm{FT}}$ as the finetuning objective, and let $\mathcal{D}_{\textrm{FT}}\subset\mathcal{D}_{\textrm{forget}}$ represent the malicious finetuning dataset, which is designed to intentionally make the model relearn concepts from the forget set. The finetuned model parameters $\theta^{\textrm{FT}}$ are updated by one or more gradient descents. For example, when using one gradient update from $\theta$, there is $\theta^{\textrm{FT}}\leftarrow\theta-\tau\cdot\nabla_{\theta}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$, where $\tau$ is the step size. The parameters $\theta$ is trained by the meta objective $\mathcal{L}_{\textrm{meta}}$ as:
$\begin{split}&\min_{\theta}\mathcal{L}_{\textrm{meta}}(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})\\ =&\mathcal{L}_{\textrm{meta}}(\theta-\tau\cdot\nabla_{\theta}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right);\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})\textrm{.}\end{split}$ (2)
To optimize the RHS of Eq. (2), the gradients are back-propagated through both $\theta$ and $\nabla_{\theta}\mathcal{L}_{\textrm{FT}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ that can be efficiently computed by automatic differentiation [41].
@@ 4.2 Meta objective $\mathcal{L}_{\textrm{meta}}$
$\begin{split}\min_{\theta}\mathcal{L}_{\textrm{meta}}&(\theta^{\textrm{FT}};\mathcal{D}_{\text{FT}},\mathcal{D}_{\text{retain}})\triangleq-\mathcal{L}_{\textrm{DM}}(\theta^{\textrm{FT}};\mathcal{D}_{\textrm{FT}})\\ &-\zeta\cdot\left[\mathcal{L}_{\textrm{DM}}(\theta^{\textrm{FT}};\mathcal{D}_{\textrm{retain}})-\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{retain}})\right]\textrm{,}\!\!\!\!\end{split}$ (3)
@@
$\begin{split}\mathcal{L}_{\textrm{meta}}\approx&-\mathcal{L}_{\textrm{DM}}(\theta;\mathcal{D}_{\textrm{FT}})+\tau\cdot{\color[rgb]{1.0,0.44,0.37}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.44,0.37}\left\|\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)\right\|_{2}^{2}}\\ &+\tau\zeta\cdot{\color[rgb]{1.0,0.44,0.37}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.44,0.37}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)}\textrm{,}\end{split}$ (4)
@@
Figure 4: Images generated by harmful prompts. The top panel displays images generated using the original SDXL model for harmful prompts. Following panels show the images generated using unlearned and meta-unlearned SDXL models before finetuning (FT), after FT on the HRM-m dataset for 100 steps, and after FT on the CLEAN dataset for 100 steps, respectively. The left three columns display images generated by ESD-u-1 and its meta variant, while the right three columns display images generated by UCE and its meta variant.
@@
Remark. As in Fig. 1, minimizing $\left\|\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)\right\|_{2}^{2}$ decreases gradient norm and thus delays relearning of forget set. Minimizing $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$ induces a $90^{\circ}$ angle between the gradients $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ and $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$, such that when the model is finetuned along $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ (the loss $\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)$ decreases), the knowledge inside the retain set will self-destruct (namely, the loss $\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$ increases).
[…]
5 Experiments
Below we first describe the basic setups of our experiments:
Base models. We choose SD-v1-4 [51] and SDXL [47] as the base models for their widespread use.
Datasets. We use SD-v1-4 and SDXL to generate both $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{retain}}$ for meta-unlearning. Subsequently, we employ FLUX.1 [26] to create three safety-related finetuning datasets HRM-s, HRM-m, CLEAN for evaluation, by applying a single harmful prompt, multiple harmful prompts and benign prompts, respectively. We also utilize FLUX.1 to create two finetuning datasets related to copyright and two others related to image style. Detailed information can be found in Appendix E.1. Additionally, the 10K subset of COCO-30K [30] is used to evaluate the generation quality of unlearned DMs while the nudity subset in the Inappropriate Image Prompts (I2P) dataset [54] is used to test the unlearning performance.
Baselines. We use four established unlearning methods as baselines, including ESD [6] and SDD [24], which remove the target concept through gradual optimization; UCE [7] and RECE [10] that achieve target concept erasure through closed-form solutions. Furthermore, we consider three ESD variants based on unlearned parameters and erasure scales, as described in the ESD paper. We use the ESD-u-1, which erases U-Net models excluding cross-attention parameters under weak erase scale $\eta=1$, ESD-u-3, which erases the same parameters as ESD-u-1 but under strong erase scale $\eta=3$, and ESD-f-3, which erases the full parameters of the U-Net with erase scale $\eta=3$.
Evaluation metrics. We use FID [16] and CLIP scores [15] to evaluate the model’s generation quality. To evaluate the effect of unlearning harmful content and resistance to malicious finetuning, we calculate nudity score [54] based on the percentage of nude images in all generated images.
Evaluation details. We first use an unlearned model to generate images based on COCO 10K subset, and compute the FID and CLIP scores using the generated images and COCO subset data. Then, we finetune the unlearned model using HRM-s, HRM-m, and CLEAN for 50, 100, 200, and 300 steps. Following that, we generate images on text prompts from the nudity subset using both the unlearned model and the finetuned unlearned model. Finally, we use the nudity detector [73, 54] to determine the nudity score for the generated images.
5.1 Harmful content removal
Tables 1 and 2 show the evaluation results of the unlearned and meta-unlearned SD-v1-4. Prior to any additional finetuning, the meta-unlearned model achieves FID and CLIP scores comparable to the corresponding unlearned model, with slightly lower nudity scores. After malicious finetuning on the HRM-s and HRM-m datasets, the unlearned model shows a rapid increase in nudity scores. In contrast, meta-unlearned SD yields significantly lower nudity scores than the unlearned model. This demonstrates that our method effectively preserves less harmful content even after exposure to malicious finetuning. Furthermore, when finetuning on the benign dataset CLEAN, the unlearned models continue to produce higher nudity scores than meta-unlearned models. Fig. 10 (in Appendix) shows images generated by unlearned and meta-unlearned models on benign prompts before finetuning, indicating that meta-unlearned models can produce comparable images with corresponding unlearned models. Then we show the images generated by harmful prompts in Fig. 11 (in Appendix). The unlearned and meta-unlearned models are finetuned on the HRM-m dataset for 50, 100, 200, and 300 steps. As finetuning steps increase, unlearned models rapidly relearn the ability to generate harmful images. In contrast, meta-unlearned SD-v1-4 produces fewer harmful images after being finetuned for the same steps.
Figure 6: The value of $\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{FT}}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\textrm{DM}}\left(\theta;\mathcal{D}_{\textrm{retain}}\right)$} for each step of meta-unlearning. Because the value is noisy, we use the regression line to represent a smoothed trend.
Fig. 8 (in Appendix) presents images generated on benign prompts by unlearned and meta-unlearned SDXL. As seen, for benign prompts, the meta-unlearned SDXL also achieves a high generation quality comparable to that of the corresponding unlearned method. Fig. 4 displays the harmful images generated by the unlearned and meta-unlearned models before and after being finetuned on the HRM-m and CLEAN datasets. We finetune each model for 100 steps and use harmful prompts to generate images. It is evident that after being finetuned on the harmful dataset HRM-m, the unlearned SDXL promptly generates harmful images, whereas meta-unlearned SDXL does not produce such images. Besides, after being finetuned on the benign dataset CLEAN, the unlearned models still have a probability of generating harmful images, while meta-unlearned models consistently ensure the generation of harmless images.